# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ Acknowledgements.
We derive Eq. (55). Since we have explained this derivation in Ref. [NA-FloquetSHE, ], we here describe its main points. Substituting Eq. (42) or (43) into Eq. (52), we obtain
@@
[…]
@@
[…]
@@
[…]
$=\frac{1}{\omega N}\sum_{\mathbf{k}}\sum_{a,b,c,d}\sum_{\sigma,\sigma^{\prime}}\int_{-\Omega/2}^{\Omega/2}\frac{d\omega^{\prime}}{2\pi}\sum_{m,n,l,q=-\infty}^{\infty}\{[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})]_{ml}[G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}+\omega)]_{ln}[v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})]_{nq}[G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]_{qm}$
@@
Since $\sigma_{yx}^{\textrm{Q}(1)}(\omega)$ has only a pure imaginary part Eckstein , we have
@@
$=$ $\frac{1}{2\omega}\Bigl{\{}\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}+\omega)v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]-\textrm{tr}[G_{a\sigma d\sigma^{\prime}}^{<}(\mathbf{k},\omega^{\prime})v_{dc\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{c\sigma^{\prime}b\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}+\omega)v_{ba\sigma}^{(\textrm{Q})y}(\mathbf{k})]\Bigr{\}},$ (122)
@@
$=$ $\frac{1}{2\omega}\Bigl{\{}\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})G_{b\sigma c\sigma^{\prime}}^{<}(\mathbf{k},\omega^{\prime})v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}-\omega)]-\textrm{tr}[G_{a\sigma d\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}-\omega)v_{dc\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{c\sigma^{\prime}b\sigma}^{<}(\mathbf{k},\omega^{\prime})v_{ba\sigma}^{(\textrm{Q})y}(\mathbf{k})]\Bigr{\}}.$ (123)
@@
Appendix B $[\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
We can calculate $[\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL by using Eqs. (12) and (13) and Eq. (83). To do this, we use identities,
@@ Appendix B $[\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{l^{\prime\prime}}-e^{-ik_{x}}e^{ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}-e^{ik_{x}}e^{-ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}],$ (129)
@@
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{l^{\prime\prime}}+e^{-ik_{x}}e^{ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}+e^{ik_{x}}e^{-ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}],$ (130)
@@
As described in Sec. III, $t_{1}$, $t_{2}$, and $t_{3}$ are the nearest-neighbor hopping integrals on the square lattice, and $t_{4}$ and $t_{5}$ are the next-nearest-neighbor ones NA-FloquetSHE [Fig. 5(a)].
We make three comments about Eqs. (127)–(130). First, the relative phase difference in BCPL, $\theta$, causes the phase factors, such as $e^{-il\theta}$ and $e^{-i(l^{\prime}+l^{\prime\prime})\theta}$, only for the terms of the finite-order Bessel functions. Second, the number of the Bessel functions is twice that in the case of CPL NA-FloquetSHE . This is because the $x$ or $y$ component of the pump field contains two trigonometric functions in the case of BCPL, whereas it contains one in the case of CPL NA-FloquetSHE . Third, even if the Floquet indices $m$, $n$, $l$, $l^{\prime}$, and $l^{\prime\prime}$ are restricted to the range of $-n_{\textrm{max}}\leq m,n,l,l^{\prime},l^{\prime\prime}\leq n_{\textrm{max}}$, $[\epsilon_{ab}(\mathbf{k})]_{mn}$’s include not only the Bessel functions the order of which is within this range, but also the Bessel functions the order of which is outside of it. For example, if $\beta=2$ and $-1\leq m,n,l\leq 1$ (i.e., $n_{\textrm{max}}=1$), the order of the first Bessel function in Eq. (127), $\mathcal{J}_{n-m-\beta l}(u)$, takes the value outside of that range. Since the third property is also related to the number of the trigonometric functions appearing in the Peierls phase factors, this suggests that the high-order Bessel functions play a more important role in the case of BCPL than in that of CPL. In fact, this is consistent with the difference between the $n_{\textrm{max}}$ dependences of $\sigma_{yx}^{\textrm{S}}$ obtained for BCPL and CPL.
Appendix C $[\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
We can calculate $[\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL in a similar way to the calculation explained in Appendix B. After such calculation using Eqs. (15), (16), and (95), we obtain
@@ Appendix C $[\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}-e^{-ik_{x}}e^{ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}-e^{ik_{x}}e^{-ik_{y}}(-1)^{l^{\prime\prime}}],$ (134)
@@
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}+e^{-ik_{x}}e^{ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}+e^{ik_{x}}e^{-ik_{y}}(-1)^{l^{\prime\prime}}].$ (135)
These equations and Eqs. (127)–(130) show that the differences between $[\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$’s and $[\epsilon_{ab}(\mathbf{k})]_{mn}$’s are the differences in the phase factors due to $\theta$ and the sign factors such as $(-1)^{l}$ and $(-1)^{n-m-\beta l}$. These differences are reasonable because $\mathbf{A}_{\textrm{BCPL}}(-t)$ and $\mathbf{A}_{\textrm{BCPL}}(t)$ are given by Eqs. (15) and (16) and by Eqs. (12) and (13), respectively.
@@ Appendix D Additional numerical results
Figure 18: The $n_{\textrm{max}}$ dependences of (a) $\sigma_{yx}^{\textrm{S}}$ and (b) $\sigma_{yx}^{\textrm{C}}$ as functions of $u(=eA_{0})$ for Sr2RuO4 driven by BCPL at $\beta=2$, $\theta=\frac{\pi}{4}$, and $\Omega=6$ eV. Here $n_{\textrm{max}}$ is the upper limit of the summation over the Floquet indices. The values of $N_{x}=N_{y}$ and $\Delta\omega^{\prime}$ used in their numerical calculations are different from those used in the numerical calculations for the results shown in Sec. V: $N_{x}=N_{y}=64$ and $\Delta\omega^{\prime}=0.01$ eV were used in the former calculations, whereas $N_{x}=N_{y}=100$ and $\Delta\omega^{\prime}=0.005$ eV were used in the latter calculations. Meanwhile, the values of $\Gamma$ and $T_{\textrm{b}}$ are the same.
We show additional numerical results to discuss the validity of our choice of the value of $n_{\textrm{max}}$. Figures 18(a) and 18(b) show the $n_{\textrm{max}}$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$, $\theta=\frac{\pi}{4}$, and $\Omega=6$ eV. In obtaining these results, we chose some parameters to be different from the values used in the numerical results shown in Sec. V: we set $N_{x}=N_{y}=64$ and $\Delta\omega^{\prime}=0.01$ eV to reduce the cost of the numerical calculations. These figures show that the results obtained for $n_{\textrm{max}}=1$ and $2$ are qualitatively the same. Therefore, $n_{\textrm{max}}=1$ may be reasonable to study qualitative properties of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\Omega=6$ eV.

### v2

@@ Acknowledgements.
We derive Eq. (55). Since we have explained this derivation in Ref. [[11]], we here describe its main points. Substituting Eq. (42) or (43) into Eq. (52), we obtain
@@
$\sigma_{yx}^{\textrm{Q}(1)}(\omega)=\int_{0}^{T_{\textrm{p}}}\frac{dt_{\textrm{av}}}{T_{\textrm{p}}}\int_{-\infty}^{\infty}dt_{\textrm{rel}}e^{i\omega t_{\textrm{rel}}}\sigma_{yx}^{\textrm{Q}(1)}(t_{\textrm{av}}+\frac{t_{\textrm{rel}}}{2},t_{\textrm{av}}-\frac{t_{\textrm{rel}}}{2}),$ (107)
$\sigma_{yx}^{\textrm{Q}(2)}(\omega)=\int_{0}^{T_{\textrm{p}}}\frac{dt_{\textrm{av}}}{T_{\textrm{p}}}\int_{-\infty}^{\infty}dt_{\textrm{rel}}e^{i\omega t_{\textrm{rel}}}\sigma_{yx}^{\textrm{Q}(2)}(t_{\textrm{av}}+\frac{t_{\textrm{rel}}}{2},t_{\textrm{av}}-\frac{t_{\textrm{rel}}}{2}).$ (108)
We calculate the right-hand sides of Eqs. (107) and (108) using Eqs. (44) and (50), the Floquet representation of the Green’s functions, and two relations. One of the two relations is that the Floquet representation of a function $A(t,t^{\prime})=\int dt^{\prime\prime}B(t,t^{\prime\prime})C(t^{\prime\prime},t^{\prime})$ is given by
@@
$\sigma_{yx}^{\textrm{Q}(1)}(\omega)$ $=\int_{0}^{T_{\textrm{p}}}\frac{dt_{\textrm{av}}}{T_{\textrm{p}}}\int_{-\infty}^{\infty}dt_{\textrm{rel}}e^{i\omega t_{\textrm{rel}}}\sigma_{yx}^{\textrm{Q}(1)}(t_{\textrm{av}}+\frac{t_{\textrm{rel}}}{2},t_{\textrm{av}}-\frac{t_{\textrm{rel}}}{2})$
@@
$\sigma_{yx}^{\textrm{Q}(2)}(\omega)$ $=\int_{0}^{T_{\textrm{p}}}\frac{dt_{\textrm{av}}}{T_{\textrm{p}}}\int_{-\infty}^{\infty}dt_{\textrm{rel}}e^{i\omega t_{\textrm{rel}}}\sigma_{yx}^{\textrm{Q}(2)}(t_{\textrm{av}}+\frac{t_{\textrm{rel}}}{2},t_{\textrm{av}}-\frac{t_{\textrm{rel}}}{2})$
$=\frac{1}{\omega N}\sum_{\mathbf{k}}\sum_{a,b,c,d}\sum_{\sigma,\sigma^{\prime}}\int_{-\Omega/2}^{\Omega/2}\frac{d\omega^{\prime}}{2\pi}\sum_{m,n,l,q=-\infty}^{\infty}$
$\times\{[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})]_{ml}[G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}+\omega)]_{ln}[v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})]_{nq}[G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]_{qm}$
@@
Since $\sigma_{yx}^{\textrm{Q}(1)}(\omega)$ has only a pure imaginary part [49], we have
@@
$=$ $\frac{1}{2\omega}\Bigl{\{}\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}+\omega)v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
$\ \ \ \ \ -\textrm{tr}[G_{a\sigma d\sigma^{\prime}}^{<}(\mathbf{k},\omega^{\prime})v_{dc\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{c\sigma^{\prime}b\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}+\omega)v_{ba\sigma}^{(\textrm{Q})y}(\mathbf{k})]\Bigr{\}},$ (122)
@@
$=$ $\frac{1}{2\omega}\Bigl{\{}\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})G_{b\sigma c\sigma^{\prime}}^{<}(\mathbf{k},\omega^{\prime})v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}-\omega)]$
$\ \ \ \ \ -\textrm{tr}[G_{a\sigma d\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}-\omega)v_{dc\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{c\sigma^{\prime}b\sigma}^{<}(\mathbf{k},\omega^{\prime})v_{ba\sigma}^{(\textrm{Q})y}(\mathbf{k})]\Bigr{\}}.$ (123)
@@
Appendix B $\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
We can calculate $\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL by using Eqs. (12) and (13) and Eq. (83). To do this, we use identities,
@@ Appendix B $\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{l^{\prime\prime}}$
$-e^{-ik_{x}}e^{ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}-e^{ik_{x}}e^{-ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}],$ (129)
@@
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{l^{\prime\prime}}+e^{-ik_{x}}e^{ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}$
$+e^{ik_{x}}e^{-ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}],$ (130)
@@
As described in Sect. 3, $t_{1}$, $t_{2}$, and $t_{3}$ are the nearest-neighbor hopping integrals on the square lattice, and $t_{4}$ and $t_{5}$ are the next-nearest-neighbor ones [11] [Fig. 5(a)].
We make three comments about Eqs. (127)–(130). First, the relative phase difference in BCPL, $\theta$, causes the phase factors, such as $e^{-il\theta}$ and $e^{-i(l^{\prime}+l^{\prime\prime})\theta}$, only for the terms of the finite-order Bessel functions. Second, the number of the Bessel functions is twice that in the case of CPL [11]. This is because the $x$ or $y$ component of the pump field contains two trigonometric functions in the case of BCPL, whereas it contains one in the case of CPL [11]. Third, even if the Floquet indices $m$, $n$, $l$, $l^{\prime}$, and $l^{\prime\prime}$ are restricted to the range of $n_{\textrm{max}}\leq m,n,l,l^{\prime},l^{\prime\prime}\leq n_{\textrm{max}}$, $\epsilon_{ab}(\mathbf{k})]_{mn}$’s include not only the Bessel functions the order of which is within this range, but also the Bessel functions the order of which is outside of it. For example, if $\beta=2$ and $1\leq m,n,l\leq 1$ (i.e., $n_{\textrm{max}}=1$), the order of the first Bessel function in Eq. (127), $\mathcal{J}_{n-m-\beta l}(u)$, takes the value outside of that range. Since the third property is also related to the number of the trigonometric functions appearing in the Peierls phase factors, this suggests that the high-order Bessel functions play a more important role in the case of BCPL than in that of CPL. In fact, this is consistent with the difference between the $n_{\textrm{max}}$ dependences of $\sigma_{yx}^{\textrm{S}}$ obtained for BCPL and CPL.
Appendix C $\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
We can calculate $\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL in a similar way to the calculation explained in Appendix B. After such calculation using Eqs. (15), (16), and (95), we obtain
@@ Appendix C $\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}-e^{-ik_{x}}e^{ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}$
$-e^{ik_{x}}e^{-ik_{y}}(-1)^{l^{\prime\prime}}],$ (134)
@@
$\times$ $[e^{-ik_{x}}e^{-ik_{y}}(-1)^{l+l^{\prime}+l^{\prime\prime}}+e^{ik_{x}}e^{ik_{y}}(-1)^{n-m-l-(l^{\prime}+l^{\prime\prime})\beta}+e^{-ik_{x}}e^{ik_{y}}(-1)^{n-m-(l^{\prime}+l^{\prime\prime})\beta+l^{\prime}}$
$+e^{ik_{x}}e^{-ik_{y}}(-1)^{l^{\prime\prime}}].$ (135)
These equations and Eqs. (127)–(130) show that the differences between $\bar{\epsilon}_{ab}(\mathbf{k})]_{mn}$’s and $\epsilon_{ab}(\mathbf{k})]_{mn}$’s are the differences in the phase factors due to $\theta$ and the sign factors such as $(-1)^{l}$ and $(-1)^{n-m-\beta l}$. These differences are reasonable because $\mathbf{A}_{\textrm{BCPL}}(-t)$ and $\mathbf{A}_{\textrm{BCPL}}(t)$ are given by Eqs. (15) and (16) and by Eqs. (12) and (13), respectively.
@@ Appendix D Additional numerical results
We show additional numerical results to discuss the validity of our choice of the value of $n_{\textrm{max}}$. Figures 18(a) and 18(b) show the $n_{\textrm{max}}$ dependences of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\beta=2$, $\theta=\frac{\pi}{4}$, and $\Omega=6$ eV. In obtaining these results, we chose some parameters to be different from the values used in the numerical results shown in Sect. 5: we set $N_{x}=N_{y}=64$ and $\Delta\omega^{\prime}=0.01$ eV to reduce the cost of the numerical calculations. These figures show that the results obtained for $n_{\textrm{max}}=1$ and $2$ are qualitatively the same. Therefore, $n_{\textrm{max}}=1$ may be reasonable to study qualitative properties of $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$ for Sr2RuO4 driven by BCPL at $\Omega=6$ eV.
Figure 18: (Color online) The $n_{\textrm{max}}$ dependences of (a) $\sigma_{yx}^{\textrm{S}}$ and (b) $\sigma_{yx}^{\textrm{C}}$ as functions of $u(=eA_{0})$ for Sr2RuO4 driven by BCPL at $\beta=2$, $\theta=\frac{\pi}{4}$, and $\Omega=6$ eV. Here $n_{\textrm{max}}$ is the upper limit of the summation over the Floquet indices. The values of $N_{x}=N_{y}$ and $\Delta\omega^{\prime}$ used in their numerical calculations are different from those used in the numerical calculations for the results shown in Sect. 5: $N_{x}=N_{y}=64$ and $\Delta\omega^{\prime}=0.01$ eV were used in the former calculations, whereas $N_{x}=N_{y}=100$ and $\Delta\omega^{\prime}=0.005$ eV were used in the latter calculations. Meanwhile, the values of $\Gamma$ and $T_{\textrm{b}}$ are the same.